OpenStudy (anonymous):

Solve 2x + 3y=9 for y.

OpenStudy (anonymous):

I got 9-2x/3

OpenStudy (mathstudent55):

You want y by itself on the left side. First, subtract 2x from both sides. Then divide both sides by 3.

OpenStudy (mathstudent55):

\(2x + 3y = 9\) \(3y = -2x + 9 \) \(y = -\dfrac{2}{3}x + \dfrac{9}{3} \) \( y = -\dfrac{2}{3}x + 3 \)

OpenStudy (anonymous):

So I was right

OpenStudy (mathstudent55):

Yes.